### v1

# Enhancing Video Large Language Models with Structured Multi-Video Collaborative Reasoning (early version)

Abstract
Despite the prosperity of the video language model, the current pursuit of comprehensive video reasoning is thwarted by the inherent spatio-temporal incompleteness within individual videos, resulting in hallucinations and inaccuracies. A promising solution is to augment the reasoning performance with multiple related videos. However, video tokens are numerous and contain redundant information, so directly feeding the relevant video data into a large language model to enhance responses could be counterproductive. To address this challenge, we propose a multi-video collaborative framework for video language models. For efficient and flexible video representation, we establish a Video Structuring Module to represent the video’s knowledge as a spatio-temporal graph. Based on the structured video representation, we design the Graph Fusion Module to fuse the structured knowledge and valuable information from related videos into the augmented graph node tokens. Finally, we construct an elaborate multi-video structured prompt to integrate the graph, visual, and textual tokens as the input to the large language model. Extensive experiments substantiate the effectiveness of our framework, showcasing its potential as a promising avenue for advancing video language models.
Figure 1: Video question-answering pipeline under different video collaboration strategies. (a) Single-video reasoning pipeline; (b) Direct multi-video collaboration pipeline: concatenate multiple video’s visual tokens, which is burdensome; (c) Structured multi-video collaboration pipeline (ours).
Figure 2: Video question answering examples from video language models. Single-video reasoning (Left): In video-1, the environmental visual cues are hard for the model to perceive, leading to a ’sports team’ hallucination based on only the textual query and linguistic priors. In Video 2, ice-related visuals are missing, and limited bartending knowledge causes the model to skip the question. Multi-video reasoning (Right): Introducing relevant videos allows the model to complete and summarize domain-specific knowledge (such as environmental protection or bartending in this case), leading to more reliable and accurate answers.
1 Introduction
The significant advancement in Large Language Models(LLMs) gives rise to the prosperity of video language models. Equipped with LLMs’ strong general knowledge, video language models showcase an impressive potential to comprehend complex video contents [27, 57, 60, 24, 36, 28, 34]. These models effectively bridge the semantic chasm between video and language, facilitating a nuanced integration of visual perception with language processing.
However, it remains an elusive goal to achieve reliable and comprehensive video reasoning. Specifically, the spatio-temporal knowledge captured from an individual video is often incomplete due to the limited visual perception and comprehending ability. Factors hindering visual perception include sparse sampling [30, 56, 58, 38], spatial occlusion [20, 9], and perspective switching [48, 12, 42, 62]). The comprehending ability is mainly constrained by the complexity of video content [53, 18] and limited effective context length [32, 17]. Since video language models mostly originate from large language models [40, 19] with the help of visual instruction tuning [31], video language models can only resort to the language prior [13, 29] when faced with knowledge incompleteness. This tendency brings spurious hallucination and misalignment with the visual facts [26, 14] to the answers.
The mainstream of video language models follows a single-video reasoning pipeline (shown in Figure 1(a)). Motivated by the research in multi-data collaboration [46, 63, 52] and retrieval-augmented generation [22, 11, 15], a promising idea to address this limitation is integrating knowledge from multiple related videos. We illustrate some video question-answering examples in Figure 2. As the examples demonstrate, single-video reasoning suffers from incomplete and redundant information from an individual video. These challenges prevent video language models from grasping valuable knowledge and lead to hallucinations and failures in answering the given questions. Introducing multiple highly related videos within the same domain can compensate for the missing information and enable the video language models to conclude and provide reliable answers.
Despite the apparent advantages, the incremental information from multiple videos does not necessarily lead to better reasoning in real cases. As illustrated in Figure 1(b), A direct multi-video joint strategy is to concatenate the visual features of multiple videos in the prompt before passing them to the video language models, which can bring an overwhelming number of tokens to video language models [24, 28, 36]. When exposed to long context, empirical evidence [17, 32] proves that LLMs are inclined to concentrate only on specific and limited segments of the input, neglecting potentially critical information. Moreover, the high dimensionality, redundancy, and obfuscation of video content contrast sharply with the structured, clear, and rule-based nature of language. This temporal visual complexity makes the synthesis and refinement of multi-video knowledge particularly challenging.
In this paper, we present a structured multi-video collaborative reasoning framework. As illustrated in Figure 1(c), Our framework achieves the alignment and completion of multi-video information at the graph-structured representation level. Under this framework, to address the redundancy and complexity of video information, we first design the Video Structuring Module (VSM). This module leverages detection and tracking to extract key targets and their spatio-temporal relationships, thereby constructing a data-efficient structured video representation. After obtaining the structured representation, we propose the Graph Fusion Module (GFM) to yield LLM-friendly graph tokens. GFM integrates the structural information into node features through the graph attention network (GAT) and then fuses the multi-video knowledge based on cross-graph attention. Finally, we arrange the fused graph tokens, target video tokens, and text tokens with an elaborately designed prompt to augment VLMs with the multi-video structured knowledge. Overall, our contributions are three-fold:
  • We explore a feasible structured reasoning framework for multi-video collaboration on video language models.
  • Under this framework, we develop the Video Structuring Module to obtain the data-efficient spatio-temporal graph for video information. Then, we further design the Graph Fusion Module to fuse the spatio-temporal relations and cross-video knowledge into the graph tokens.
  • Extensive experiments show that our framework can boost the reliability and accuracy through multi-video structuring and collaboration.
2 Related Work
2.1 Vision Language Models
In the field of visual understanding, significant advancements have been achieved along with the development of large language models. By learning joint the representation of vision and language through feature alignment and visual instruction tuning [31], vision-language models leverage the generalist capabilities that emerge from the large-scale training of large language models, enabling understanding of images and videos in open-world settings. Among them, Video-LLaMA [60] and BLIP-2 [23] adopt the Q-former to extract valuable visual information into compact and LLM-friendly visual tokens for alignment. On the contrary, Flamingo [1] and BLIP-3 [55] utilize the scalable perceiver resampler to obtain learnable visual tokens. LLaVA [31], Video-LLaVA [28], and Qwen2-VL [43] use an MLP-based projector to map the visual and textual inputs into a shared feature space. Besides, due to the complicated information contained in image and video input, many studies [43, 23, 27] aim to reduce the number of visual tokens for alleviating the computational and understanding burden on LLMs. While current methods focus on single-video reasoning, we try to explore a structured video reasoning framework to effectively address the challenges of redundancy and integration brought by multi-video inputs.
2.2 Multi-Data Collaboration
Although most deep learning task follows a single data processing pipeline, multi-data collaboration provides a promising direction to improve performances with the inner correspondence among multiple samples. We can roughly divide current works into two streams, which are content-related and task-related collaborations. Content-related collaboration helps models to focus on vital content through comparison among several related data. For example, co-segmentation task [63, 46] explores ways to segment the same object in different scenes by concluding and sharing the same object-related features. Few-shot image classification [16], action recognition [44, 4], and fine-grained classification [35] methods focus on the key differences for accurate classification through multi-data comparison. Retrieval-Augmented Generation [22, 11, 15] is also a promising content-related collaboration that provides the LLMs with supporting information from retrieved related data. In task-related collaboration, models learn the way to complete the task by observing multiple samples performing the same task. For example, multi-video summarization [51, 52, 37] researches aim to generate a summary from a collection of videos through multi-data complementary and refinement. In-Context Learning [45, 3] uses examples with task guidance and answering to demonstrate to LLMs how to fulfill the tasks. For video language models, current multi-video collaboration strategy directly concatenates multiple inputs, leading to burdensome redundancy and difficulties in collaboration. Building on these advancements, we introduce the multi-video reasoning task and achieve the compensation and refinement of key spatio-temporal information through multi-video collaboration.
Figure 3: Multi-video collaborative reasoning framework. Together with the target video, NN related videos are retrieved to facilitate the reasoning process. First, we design the Video Structuring Module to obtain the structured video representation. Then, the Graph Fusion Module fuses the structure information and related videos’ information to get the video graph tokens. Finally, according to designed prompts, the graph tokens, visual tokens and text tokens are arranged as input to the large language model for question answering.
3 Method
3.1 Overview
3.1.1 The setting of multi-video reasoning
In this paper, we explore the potential of utilizing multi-video information for video understanding. Specifically, we follow a multi-video setting, where a target video $V_{0}$ is accompanied by $N$ related videos $\{V_{1},V_{2},\dots,V_{N}\}$. To retrieve related videos, their feature vectors are constructed in advance for efficient video retrieval. Different ways of video vectorization are discussed in supplementary materials. Finally, the methods are required to answer the question to the target video with the help of $N$ retrieved videos.
3.1.2 Framework overview
Our multi-video collaborative reasoning framework is illustrated in Figure 3. First, the Video Structuring Module is introduced in Sec. 3.2 to get the structured video representation. Based on the obtained video structure, we design the Graph Fusion Module in Sec. 3.3 to fuse the structure information and select useful related videos’ information to the resulting graph tokens. Finally, we arrange all graph, visual, and text tokens according to our designed multi-video reasoning prompt and send them to the large language model.
3.2 Video Structuring Module
Efficient structured video representation paves the way for the following integration of multi-video knowledge. Given the video and its paired dense caption, the process of our Video Structuring Module (VSM) is depicted on the right bottom of Figure 3, which can be specified as follows.
Step 1: Scene detection. To avoid the video’s temporal redundancy, we use a light-weighted content-based scene detector PySceneDetect [8] to detect scenes from the video. Then, the middle frame of each scene is extracted as the keyframes for the later video structuring process.
Step 2: Dense video captioning. To prepare for the latter structuring pipeline, we need to extract the detailed and fine-grained textual concept. To achieve this goal, we adopt a video language model to give a detailed description of the input video through prompt design, which is detailed in our supplementary materials.
Step 3: Textual scene graph parsing. We then extract the textual scene graph $\mathcal{G}^{\text{Text}}$ from the dense video caption with the help of the SceneGraphParser [50]. The textual scene graph is composed of several triplets $\tau_{i}=\{s_{i},p_{i},o_{i}\}$, where $\tau_{i}$ represents the $i$-th triplets and $s_{i},p_{i},o_{i}$ represents the subject, predicate and object of $\tau_{i}$. Triplets are fundamental objects of interaction and event within the video, formatted in ⟨subject - predicate - object⟩.
Step 4: Graph-guided grounding. Based on the textual scene graph $\mathcal{G}^{\text{Text}}$, we can detect the interacted objects using the open-set detector GroundingDINO [33]. Then, we use the tracking method DEVA [5] to establish the tracklets of the former grounded objects, which temporally connect the same objects. Through detecting and tracking, we can get the: 1) grounding results $\mathcal{B}=\{(b_{s_{i}},b_{o_{i}})|1\leq i\leq M$}, where $b_{s_{i}},b_{o_{i}}$ are the bounding boxes of $s_{i}$ and $o_{i}$, and 2) tracking results $\mathcal{T}=\{\text{Tid}_{s_{i}},\text{Tid}_{o_{i}}\}$, where $\text{Tid}_{s_{i}},\text{Tid}_{o_{i}}$ are the tracking ID of $s_{i}$ and $o_{i}$. Notice that triplets without grounded objects are redundant and should be filtered.
Step 5: Video graph establishment. Built upon $\mathcal{B}$, $\mathcal{T}$ and $\mathcal{G}^{\text{Text}}$, we establish the graph representation for both the target video and related videos. Our graph consists of the nodes, intra-frame edges and inter-frame edges. The node represents the features of the object in the video. To get the node-level features, we use the detecting results $\mathcal{B}$ to crop them from the keyframes and use OpenCLIP [6] to extract their visual features. Specifically, we only utilize the [CLS] token as the output. Since the regions are pre-cropped, the [CLS] token already retains a fine-grained feature representation, eliminating the need to use all output tokens. This significantly reduces the number of tokens required to represent the visual modality. The intra-frame edge is represented by the predicates of each triplet, which indicates the spatial and interacting relationships between objects within one frame. The inter-frame edge connects objects in different frames with the same tracking ID, indicating their temporal relationships. Through the above steps, we can establish the graph-structured representation of both the target video and its related videos for further collaboration.
3.3 Graph Fusion Module
The Graph Fusion Module (GFM) is a multi-layer stacked architecture (see Figure 3), with each layer comprising two key components: the Hierarchical Frame Graph Attention Network (HF-GAT) [41] and the Cross-Graph Attention (CGA) mechanism.
HF-GAT is tailored to fuse graph-structured data within a single video. Traditional Graph Attention Networks (GATs) are primarily used for tasks such as node classification and relation prediction, where relationships between nodes are explicitly defined. In contrast, for both inter-frame and intra-frame contexts, these relationships are often implicit or absent. To address this challenge, as introduced in Section 3.2, we propose converting the original visual modality data into a graph-structured representation. In this graph, nodes represent features of subjects or objects, which are extracted using OpenCLIP [6]. For intra-frame edges, we utilize relationships based on triplets, where a directional link is formed from $s_{i}$ to $o_{i}$. For inter-frame edges, we leverage the results from the tracker $\mathcal{T}$ to link $s_{i}^{t-1}$ or $o_{i}^{t-1}$ from the previous frame to $s_{i}^{t}$ or $o_{i}^{t}$ in the current frame. Additionally, we introduce bidirectional links for inter-frame connections, as it enhances the system’s ability to comprehend video content [39].
Once the structured features from individual videos are extracted using HF-GAT, the next step is to select the most relevant information from related videos for the target video. To accomplish this, we introduce the cross-graph attention mechanism. Specifically, we use the target video $\mathbf{X}_{tar}$ as the query $\mathbf{Q}_{tar}=\mathbf{X}_{tar}\mathbf{W}_{Q}$, while the concatenation of the target video and related videos along the frame dimension serves as the key and value, i.e. $\mathbf{V}_{tar},\mathbf{V}_{rel}]=[\mathbf{X}_{tar},\mathbf{X}_{rel}]\mathbf{W}_{V}$ and $\mathbf{K}_{tar},\mathbf{K}_{rel}]=[\mathbf{X}_{tar},\mathbf{X}_{rel}]\mathbf{W}_{K}$, where $[,]$ denotes concatenation operation along the frame dimension. The concatenation order of related videos is unordered and interchangeable, so we do not explicitly introduce positional encoding, drawing inspiration from the Set Transformer [21]. The positional encoding within individual videos is implicitly provided by the structured information embedded in the data, as the input to the cross-graph attention mechanism is derived from the output of HF-GAT.
Furthermore, we use class embedding (CE) to enable the cross-graph attention mechanism to distinguish between the graph of the target video and the graphs of related videos.
$\mathrm{CE}_{tar}$ $=\sigma(\boldsymbol{\alpha}),$ (1)
$\mathrm{CE}_{rel}$ $=1-\sigma(\boldsymbol{\alpha}),$
where $\boldsymbol{\alpha}$ represents a learnable parameter shared across frames. We then apply CE directly to the query, key, and value by adding it to each component:
$\mathbf{\hat{Q}}_{tar}=\mathbf{Q}_{tar}+\mathrm{CE}_{tar},$ (2)
$\mathbf{\hat{K}}_{rel}=\mathbf{K}_{rel}+\mathrm{CE}_{rel},\quad\mathbf{\hat{V}}_{rel}=\mathbf{V}_{rel}+\mathrm{CE}_{rel}.$
Figure 4: Structured multi-video prompts. We properly integrate the multi-modal tokens, together with the one-shot and Chain-of-Thought guidances, to form an LLM-friendly input.
By leveraging CE, the cross-graph attention mechanism implicitly learns to distinguish between the target video and the related videos. The final output of the cross-graph attention is produced by concatenating $\mathbf{O}_{tar}$, the fused representation of the target video, with $\mathbf{X}_{rel}$, the original representation of the related videos that are initially fed into the attention mechanism.
We apply residual connections and pre-norm for both the HF-GAT and cross-graph attention components. Specifically, we use RMSNorm [40], which is commonly employed in large language models (LLMs), for pre-norm. Notably, we omit the Feed-Forward Network (FFN) in this layer to maintain the alignment of features without excessive shifting. The graph-structured representations of the videos are then processed through the GFM to construct graph tokens, each corresponding to the features of subjects or objects after being fused with the structured information.
3.4 Structured Multi-Video Prompt
Upon obtaining the fused multi-video graph tokens, we need to integrate the graph, video, and text tokens together to create an LLM-friendly prompt. Therefore, we propose the structured multi-video prompt, as is shown in Figure 4. Our prompt originates from previous video language models’ prompt design [28]. For target video, we maintain the video tokens <<VIDEO_TOKENS>> of the target video to preserve fine-grained and detailed visual information. We also append its graph-structured data <<GRAPH_TOKENS>> to indicate the key object and spatio-temporal relationships. For $N$ related videos, we only maintain the concise and data-efficient graph-structured data <<GRAPH_TOKENS>>. In this context, we further indicate the relationships between the target video and related videos and how the LLMs can utilize these related multi-video structured data. Moreover, we introduce Chain-of-Thought (CoT) [49] reasoning through the fixed prompt “Think through the process step by step”. This approach helps the model to sequentially reason through structured information, improving its ability to comprehend and process the graph-based features. By structuring the prompt in this manner, we enable the video language model to leverage multi-video information in an effective way, thereby enhancing the model’s ability to reason about and answer queries related to the video content.
4 Experiments
4.1 Experiment Setup
Dataset descriptions. At the training stage, we finetune our model on a subset of VideoInstruct-100K [36] dataset with 10K video-text instructions together with the structured video graph. For evaluation, we test our method on three Video-QA datasets, including two public open-ended video question-answering datasets: ActivityNet-QA [59] and MSRVTT-QA [54]. Besides, we also establish a high-quality domain-related Video-QA dataset named InternVid-QA. We meticulously selected 631 videos across 41 major domains from the large-scale video dataset InternVid [47]. We illustrated the statistics and some video examples in Figure 5. InternVid-QA’s videos cover various domains of life (see Figure 5(a)) and are annotated with high-quality, domain-relevant question-answer pairs (see Figure 5(b)). InternVid-QA’s setting aligns more closely with the objectives of our multi-video collaborative understanding. InternVid-QA will be public available for research use.
Implementation details. Our proposed structured multi-video collaboration framework is adaptive to general video language models. To conduct experiments, we uses Vicuna-7B v1.5 [7] as the language model and LanguageBind [64] based on ViT-L/14 [10] as its video encoder, following Video-LLaVA [28]. For the Graph Fusion Module, the output node-level feature’s dimension is 1024. Since Video-LLaVA has pre-trained the projection layer to align the OpenCLIP’s visual features with LLM’s text tokens, we just pass the 1024-dim node feature to the projector and get the 4096-dim graph tokens. We initialize our video language model with Video-LLaVA 7B’s model weights and empower the model with the capability to learn from multi-video graphs through finetuning. In the fine-tuning process, the vision and text encoder are frozen. For the remaining network, we use the AdamW optimizer with a cosine learning rate schedule, whose initial learning rate is $2\times 10^{-5}$ and the warmup ratio is 0.03. We train our model on 8 A6000 GPUs with a total batch size of 32. For each video, we uniformly sample 8 frames and resize each frame to $224\times 224$. Each video is accompanied by 5 related videos during training and experiments are conducted on InternVid-QA without special comments.
Figure 5: The InternVid-QA dataset. (a) Statistic results of video number distribution over 41 different domains. (b) Illustration of some video examples with question and answer labels on the “painting” and “salad” domains.
Table 1: Video question answering performances on different large video-language models. Following Video-ChatGPT [36], we employ ChatGPT-Assistant (gpt-3.5-turbo) to report the accuracy and score of the open-set answering.
Method LLM Size MSRVTT-QA ActivityNet-QA InternVid-QA
Acc Score Acc Score Acc Score
Video-ChatGPT 7B 49.3 2.8 35.2 2.7 - -
LLaMA-Adapter 7B 43.8 2.7 34.2 2.7 - -
LLaMA-VID 7B 57.7 3.2 47.4 3.3 - -
MiniGPT-4 7B - - - - 25.0 2.3
MiniGPT4-Video 7B 59.7 3.3 46.3 3.4 30.2 2.5
VideoChat2 7B 54.1 3.3 49.1 3.3 47.1 3.1
Video-LLaVA 7B 59.2 3.5 45.3 3.3 45.6 3.1
[1pt/1pt] Ours 7B 61.8 +2.6 3.5 46.9 +1.6 3.4 49.5 +3.9 3.1
4.2 Zero-shot Video Question Answering
We evaluate the advanced video language models on the open-ended video question-answering tasks under a zero-shot setting. The benchmark is composed of three datasets, involving ActivityNet-QA, MSRVTT-QA, and our newly proposed InternVid-QA. For a fair evaluation of the open-ended answers, we follow Video-ChatGPT [36] to use GPT-3.5 to evaluate the accuracy (True/False) and quality score (from 0 to 5). We compare the performances of several advanced large video-language models, including Video-ChatGPT [36], LLaMA-Adapter [61], Video-LLaVA [28], LLaMA-VID [27], MiniGPT-4 [65], MiniGPT4-Video [2] and VideoChat2 [25]. These competitors all conduct zero-shot video question answering conditioned on a single video. As mentioned above, our method’s answering is conditioned on the target video and several retrieved videos from the test set. According to the experimental results in Table 1, our framework improves the baseline (Video-LLaVA) by 2.6% on MSRVTT-QA, 1.6% on ActivityNet-QA and 3.9% on InternVid-QA, which shows its capability to effectively integrate multi-video knowledge and provide more reliable answers.
Table 2: Ablation study on video structuring and multi-video fusion components on InternVid-QA.
Struct Multi-video context $L$ Acc
single video 2.1K 45.6
multi-video tokens 12.5K 28.5
multi-video captions 2.5K 46.8
[1pt/1pt] ✓ single video 2.1K 46.1
graph fusion module 2.3K 49.5
Table 3: Ablation on the design of GFM.
HF-GAT CGA FFN Acc
46.2
47.8
49.5
48.7
4.3 Ablation study
Ablation study on the proposed components. Our framework mainly consists of two parts, which are video structuring and multi-video collaboration. We conduct ablation study on our proposed components to validate their effectiveness. Experimental results are shown in Table 2. We first disable video structuring module to test different common multi-video fusion strategies. Here “multi-video tokens” means concatenate all videos’ tokens as the input to context; “multi-video captions” means sending all videos’ captions to context. By observation, “multi-video tokens” strategy introduces overwhelming tokens to the context and brings excessive understanding challenges, yielding bad performance. “multi-video captions” can slightly improve the performance (1.2%) with the help of related videos. Then, we enable the Video Structuring Module and find that a structured and clear video representation is beneficial to LLM’s understanding, improving the performance by 0.5%. Finally, by adopting multi-video graph fusion, the framework can effectively extract useful information from related videos and get 3.4% additional performance improvement, at the cost of only 0.2K overhead of token numbers.
Ablation on Graph Fusion Module design. We further conduct an ablation study on the design of our Graph Fusion Module. Our GFM is composed of two components: HG-GAT and CGA. Since we have discussed the usage of FFN in Section 3.3, we also take it into consideration and conduct ablation study on these three components. Ablation results are shown in Table 3. In the first row, the graph structure feature tokens are directly sent to the multimodal projection layer to get the graph tokens. Then in the second row, we adopt HF-GAT to convey structure information to each token, increasing the performance by 1.6%. Then, with the help of CGA, multi-video knowledge is fused into graph tokens and further improves the performance by 1.7%. The adoption of FFN doesn’t bring improvement to the reasoning accuracy, which further corroborates the conclusion in Section 3.3 above.
Figure 6: Accuracy under different video numbers.
Figure 7: Visualization of our structured multi-video collaborative reasoning. We visualize a representative video question-answering example in our structured multi-video collaboration pipeline, including the grounding and structuring results, the cross-graph attention map, as well as the answers before and after adopting our framework.
Figure 8: Accuracy under different video relevance.
4.4 Discussion on the retrieved video contents
How do multiple videos affect the performance? Multi-video data can help to achieve more comprehensive reasoning, In this section, we discuss the effect of the number of the related videos. As illustrated in Figure 6, when the retrieval number increased from 0 to 7, we can observe that the accuracy first rises then drops, which peaks at 5.
How does video relevance affect the performance? In the main experiment, we arrange most related videos to the target video within each iteration. In this part, we rearrange the retrieved videos with different lower similarities (cosine similarity between video features) to test how video relevance affects the performance. As illustrated in Figure 8, the reasoning performance decreases as the relevance of the video diminishes, while the performance is still comparable to our competitors.
conclusion. More relevant videos can lead to better performance, but the key is to find videos containing critical information. Introducing more or unrelated videos will unavoidably introduce irrelevant information, yet our method possesses the ability to filter out irrelevant information to some extents.
4.5 Visualization
We visualize the reasoning process of our multi-video collaboration framework in Figure 7. The query is “What was the exact function of the machine used by the man?”, which requires high-level and domain-related knowledge and the baseline model failed to provide a satisfying response. In our framework, the video is represented as graph-structured data, maintaining the key spatio-temporal information. Then, through the cross-graph attention shown in Figure 7, important sub-graph structures from related videos are fused into current graph features, supporting our model to provide an accurate answer.
5 Conclusion
This paper introduces a pioneering framework that enhances video large language models through structured multi-video collaborative reasoning. We first design the Video Structuring Module that models video as a spatio-temporal graph. Then, The Graph Fusion Module integrates related video information into enhanced graph tokens, which are then combined with visual and textual tokens in a multi-video structured prompt for input to the language model. Extensive experiments demonstrate the effectiveness and robustness of our approach to comprehending complex video content and answering queries accurately. We hope our work can provide insights for reliable video understanding and stimulate more research interest.